### v1

# TreeCat: Standalone Catalog Engine for Large Data Systems

Abstract.
With ever increasing volume and heterogeneity of data, advent of new specialized compute engines, and demand for complex use cases, large scale data systems require a performant catalog system that can satisfy diverse needs. We argue that existing solutions, including recent lakehouse storage formats, have fundamental limitations and that there is a strong motivation for a specialized database engine, dedicated to serve as the catalog. We present the design and implementation of TreeCat, a database engine that features a hierarchical data model with a path-based query language, a storage format optimized for efficient range queries and versioning, and a correlated scan operation that enables fast query execution. A key performance challenge is supporting concurrent read and write operations from many different clients while providing strict consistency guarantees. To this end, we present a novel MVOCC (multi-versioned optimistic concurrency control) protocol that guarantees serializable isolation. We conduct a comprehensive experimental evaluation comparing our concurrency control scheme with prior techniques, and evaluating our overall system against Hive Metastore, Delta Lake, and Iceberg.
PVLDB Reference Format:
PVLDB, 19(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/umddb/treecat.
1. Introduction
The catalog is an important component of a database management system, responsible for storing and maintaining metadata. It contains various types of information, including the properties of logical entities such as databases, tables, views, triggers, and schemas; mapping of these entities to the underlying physical storage; and access control information, to name a few. Because metadata operations are in the hot path of most database operations, a database management system cannot operate efficiently without a high-performance catalog.
In a monolithic data system, the catalog is implemented as a component tightly coupled with the rest of the system. This design suffices when there are only one or a few data systems under operation. But, in a typical enterprise setting, this is rarely the case. Since no single one-size-fits-all system can meet the requirements of all different use cases, organizations often use many specialized engines, each serving a different function. Because each system maintains its own catalog with limited external access, data cannot be easily shared and ETL (extract-transform-load) pipelines need to be set up to migrate data between different systems. However, with the increasing variety and volume of data, along with the continuing emergence of new systems, maintaining and building these pipelines becomes increasingly cumbersome.
At the same time, there has been a growing trend toward the disaggregation of large-scale data systems into smaller subsystems. This includes the separation of compute and storage layers, the increasing adoption of common table formats (parquet, ORC), and the use of “connectors” to link different systems. These efforts aim to achieve a composable data stack, enabling organizations to store data in a shared storage layer and choose from multiple compute engines. However, as long as each system’s metadata about the “managed” data remains separate, maintaining numerous systems will continue to be challenging. For data management systems to be truly composable, the catalog functionality has to be disaggregated from the rest of the system. This issue is now widely recognized in the data systems industry, and there are multiple active projects, including lakehouse storage formats (Foundation, 2024d, c, a), and, most recently, catalog services (Databricks, 2025, 2024; Sakdeo, 2024; Snowflake, 2024a), mainly driven by the industry. However, as we elaborate in the next section, these systems have architectural designs that impose fundamental limitations on both current and future use cases.
In this context, the catalog itself can be seen as an independent database (sub-)system specialized for handling metadata operations requested by various database engines. As such, the standalone catalog has its own set of functional and systemic requirements, which are not adequately addressed by existing solutions.
In this paper, we first characterize the key requirements for an ideal standalone catalog system. We then present our prototype catalog engine, TreeCat, with the following design choices. (1) Given the hierarchical nature of metadata logical modeling and query access patterns, we adopt a hierarchical data model along with a path navigation query language similar to XPath. (2) For efficient query execution, we employ a storage layout inspired by file systems, implemented using write-optimized key-value stores, and leverage batch correlated scan operations. (3) We use versioned storage to support MVCC (multi-version concurrency control) for improved read performance, as well as version control operations such as time travel queries, clone, and snapshot. (4) We present a novel MVOCC mechanism that combines scan range locking and precision locking techniques. This approach efficiently manages predicate dependencies while minimizing validation costs. Additionally, we adopt a commit-time update technique to avoid aborts caused by conflicts on frequently updated fields, such as statistics. Although we focus on catalogs in this paper, our techniques are more generally applicable in any hierarchical database management system. TreeCat is implemented in about 12000 lines of C++ code, and exposes a gPRC interface to serve remote client requests. We present a comprehensive set of experiments illustrating the significant performance benefits of our new MVOCC protocol. We also compare TreeCat with Delta Lake, HMS, and Iceberg to show that TreeCat supports read operations and commit protocol with significantly lower latency.
Figure 1. Disaggregating the catalog functionality for higher composability
2. Background and Prior Work11footnotemark: 1
In traditional RDBMSs and most OLTP systems today, catalog metadata is stored as a predefined set of relations (Hellerstein et al., 2007; PostgreSQL, [n.d.]; MySQL, 2016). Apart from a specialized in-memory cache for low-latency reads, all metadata, including information about namespaces, databases, index, access control, and statistics, use the same underlying storage representation and concurrency control mechanism as the data. Because both metadata operations and data operations are transactional, and the volume of metadata is relatively small, such implementations are sufficient in most cases. However, the tight integration and specialized code make it difficult to extend catalog functionality beyond a single system, or to enrich it to support emerging requirements such as versioning.
With the emergence of data warehouses as primary OLAP systems, different catalog implementations have been developed to handle increasingly decentralized and federated scenarios. While implementations vary across systems, the catalog of a typical data warehouse functions as a specialized subsystem that decouples metadata from data, since the primary data processing infrastructure is not optimized for transactional workloads. Nevertheless, monolithic design persists, as the catalog remains tightly integrated with the rest of the system (Motivala, 2018; Armenatzoglou et al., 2022; Vandiver et al., 2018). For example, Snowflake, a cloud-native data warehouse, stores its metadata in FoundataionDB (Zhou et al., 2021), a distributed transactional key-value store. Many of its features, including automatic data clustering, data pruning during query execution, access control, and zero-copy clone, rely on catalog objects persisted in FoundationDB (Snowflake, 2024b; Motivala, 2018). However, because external systems cannot directly access the metadata, they cannot perform as efficiently or offer the same features when handling tables actively managed by Snowflake.
On the other hand, large-scale data processing engines (Thusoo et al., 2010; Armbrust et al., 2015; Sethi et al., 2019) from the big data community originally emerged with a highly disaggregated architecture, consisting of an independent compute engine, distributed file system, and a metadata service, each designed to be highly scalable. As part of the Hive ecosystem, HMS (Hive Metastore) (Thusoo et al., 2010) was commonly used as the standalone catalog system. Backed by a transactional RDBMS (typically MySQL or PostgreSQL), HMS stores information about databases, tables, access control information, statistics, and partitions (groups of file objects organized by common attribute values). Unfortunately, HMS has a major shortcoming: it does not maintain file-level metadata. Instead, it relies on the underlying storage system (distributed file system or object store) to group files in the same partition under the same directory, essentially using the directory structure as an index. To scan a table, a system must retrieve the paths of relevant partitions from HMS and call list() on each partition directory to obtain the list of files to scan. In environments where data could scale to millions of partitions, the high latency of list() operation of cloud object stores like S3 becomes a serious performance bottleneck (Armbrust et al., 2020). In addition, even retrieving the list of large number of partitions from HMS is slow due to an inefficient Thrift server implementation and suboptimal back-end database schema (Katz, 2021).
As these limitations of HMS became apparent, lakehouse storage formats, namely Delta Lake, Hudi, and Iceberg (Armbrust et al., 2020; Foundation, 2024a, c) emerged as alternatives. While the specific implementations differ, the common idea is to store metadata about file objects as separate metadata files along with the data files in shared storage. By directly fetching and processing these metadata files, the performance bottlenecks of HMS can be avoided. Furthermore, additional file metadata, such as statistics, can be stored in metadata files and utilized for optimizations like file pruning, further improving system performance. Delta Lake uses a data structure called Delta log, which consists of a sequence of delta JSON files (with incrementing numerical IDs) and periodic checkpoint files that store the complete list of file objects at specific snapshots. Iceberg organizes metadata into a hierarchy of files: a metadata file contains a list of manifest list files; a manifest list file contains a list of manifest files; and a manifest file contains a list of file objects. Hudi identifies files by the timestamps at which they are first added and stores a set of indices in HFile format, enabling fast pruning of files based on attributes.
Storing the metadata alongside the data provides additional benefits such as a simplified architecture, good scalability due to the underlying storage system, and speedup from parallel metadata processing. However, this approach has notable drawbacks, which are becoming increasingly problematic as lakehouse systems evolve to support more features. One major drawback is the lack of an advanced concurrency control mechanism, especially multi-table transactions. Another issue is the overhead of operating on file-level metadata, which requires fetching metadata files from shared storage before processing them, resulting in higher latency. Lastly, every lakehouse system still requires a separate catalog system to store the metadata about high-level catalog objects, such as databases and tables. More recently, several efforts have focused on providing these services through Hive-compatible APIs or REST APIs (Databricks, 2025, 2024; Sakdeo, 2024; Snowflake, 2024a). Ironically, contrary to the motivation of replacing HMS with a simpler scheme, lakehouse storage formats are becoming part of even more complex ones, where metadata spans multiple components.
Enterprise data catalogs serve as an organization-level database of metadata about datasets across various systems. There are numerous commercial data catalog software products, including those by Alation, Collibra, Informatica, AWS (Glue), Google (Dataplex), and Microsoft (Azure Data Catalog), as well as open source projects, such as Apache Atlas (Foundation, 2018), Metacat (Netflix, 2024), and DataHub (Authors, 2025a). Unlike operational catalogs, the primary goal of data catalogs is organization-wide data governance, which involves constructing and maintaining metadata stores through data discovery processes (Vahlkamp, 2024; Authors, 2025b; Majumdar and Li, 2018; Services, 2025). However, systems like AWS Glue also offer functions traditionally provided by operational catalogs. Although, the distinctions between them are becoming increasingly blurred, the performance characteristics of the enterprise data catalogs make them unsuitable for use as operational catalogs today.
3. Requirements
Some of the key requirements for a standalone operational catalog are as follows.
Functional Requirements: The key functional requirements for a catalog, as derived from our investigations, are as follows: (1) Manage metadata about numerous logical objects and their relationships, which are fundamentally hierarchical (but not necessarily tree-structured), providing semantic meaning to the underlying data. (2) Support various statistics and clustering information at the level of physical objects, which is increasingly utilized by database engines for cost-based query optimization. (3) Provide a flexible query language, or API, to support different types of queries. (4) Incorporate explicit versioning semantics to support version control operations, such as time travel queries and clone operations. (5) Be extensible, so a large variety of data sets (e.g., non-tabular data) and database engines can be supported.
System Requirements: Supporting the core functionalities described above imposes the following systems-level requirements: (1) Efficiently handle transactional read operations with low latency (fast range queries are especially important). (2) Support frequent insert, delete, and update operations that accompany ETL (extract, transform, load) processes. (3) Implement a scalable concurrency control mechanism with strong ACID guarantees.
Out-of-scope Requirements: Finally, we briefly outline important considerations that are beyond the scope of this paper, but drive our future research: (1) Garbage collection, required to remove obsolete physical files and objects, (2) Availability and fault tolerance, which is essential in a distributed architecture, and (3) User-defined Types and Functions, which are powerful features that further enhance extensibility.
In addition, there are two important issues closely aligned with the catalog, but should be done externally to the catalog. First, maintaining consistency between the data and the metadata is a major emerging challenge. However, it is unclear where that responsibility should lie. While our proposed catalog system does not enforce data-metadata consistency internally, it provides functionality for external systems to manage it. Similarly, access control is a crucial challenge in this space, but needs to be enforced by the layer that manages access to the data itself, which most catalogs (including TreeCat) don’t.
4. High Level Design
In this section, we present the high-level data model and the basic API of TreeCat.
4.1. Data Model
TreeCat adopts a hierarchical data model. A typical metadata model of a database consists of a nested hierarchy of namespaces, databases, tables, partitions, and file objects. Even in the case of non-tabular data sets, the hierarchical model is suitable because some logical grouping of the underlying data is necessary. The model takes advantage of data locality, where objects within the same logical group are often accessed together. To support various systems, the model should be structurally flexible, of arbitrary depth, and hold objects of different types, as illustrated below:
Figure 2. An example retail database, which shows a typical nested hierarchical structure of a relational database system.
Example 1: Figure 2 illustrates a retail database consisting of multiple tables, which are further partitioned by multiple attribute values. The lowest-level partitions contain physical files and database blocks.
Example 2: Figure 3 illustrates datasets containing some metrics stored in a data lake. One dataset is a data stream of active metric logs ingested in batches. The other dataset is aggregate data stored as a table and is derived from the active log data.
We now proceed with formal definitions of our data model.
  1. (1)
    Every node in the hierarchy corresponds to an object of a type with a set of properties.
  2. (2)
    There are 2 types of objects:
    1. (a)
      A non-leaf object has a parent and a set of children objects. It is uniquely identified by its path, $/oid_{1}/oid_{2}/../oid_{n}$, where $oid_{n}$ is its object id and the object with the path of $/oid_{1}/oid_{2}/.../oid_{n-1}$ is its parent.
    2. (b)
      A leaf object has one or more parents, but does not have any child object. A leaf object with multiple parents has multiple paths. Given one of its paths, $/oid_{1}/oid_{2}/../oid_{n}$, $oid_{n}$ is one of its object ids and the object with the path of $/oid_{1}/oid_{2}/.../oid_{n-1}$ is one of its parents. Unlike non-leaf objects, a leaf object is immutable and can only be added or removed.
  3. (3)
    Every change (object addition, removal, or update) is associated with a $vid$ (version id) that determines its globally consistent chronological order.
Some of our design choices deviate from standard practices, so are worth documenting:
  1. (1)
    Because leaf objects often represent physical objects (e.g., Parquet files) that are short-lived and almost never modified, they are immutable, avoiding the storage of past versions of the metadata.
  2. (2)
    Because a leaf object may have multiple non-leaf parents, the hierarchical structure of the data model takes the form of a DAG, which enables sharing datasets without duplicating the underlying physical objects or their metadata.
  3. (3)
    TreeCat provides versioning semantics due to its high utility. But, considering practical functionality, we only provide snapshot operation and clone operation, as described below.
Figure 3. An example metrics data sets, consisting of a data stream and a table.
4.2. Basic API and Query Language
Next, we list a basic API that TreeCat supports.
  • startTransaction(mode): Starts a new transaction in the specified mode. If the mode is read-only, it returns the latest $vid$ at which a snapshot of the catalog can be read. If the mode is read-write, it creates a new transaction session and returns $transaction\_id$ and $read\_vid$.
  • executeQuery(query, optional transaction_id, optional vid)): Executes a read query on the catalog using the query language, defined below.
  • commit(transaction_id, write_set): Validates and commits the transaction of identified by $transaction\_id$, applying modifications specified by the given $write\_set$.
  • snapshot(snapshot_name, vid): Creates an immutable snapshot of the catalog data at $vid$ and names it $snapshot\_name$.
  • clone(src_path, dest_path, optional vid): Clones/copies the object at $src\_path$ and all of its descendants to the $dest\_path$. Note that this is a metadata-level operation, and does not involve cloning the physical objects.
Query Language: TreeCat’s query language specifies a hierarchical path navigation. It is similar to XPath (Clark et al., 1999), but is much more constrained as it only allows top-down navigation of each context object’s children whereas XPath allows navigation along one of multiple different axes (children, descendants, following, parent etc.) at each step. A query takes form of a path expression, a sequence of $n$ predicates, each of which is evaluated against the children of context objects that satisfy the preceding path subexpression. A predicate is either a wildcard character or an expression (enclosed with [][][ ]) that may be evaluated against the object id or object properties. We extend the example in Figure 2 to construct an example query. A query that asks for files and database blocks containing sales data for clothes in Asia can be constructed as follows:
/[obj_id = "retail"]/[name = "Sales"]/[region = "Asia"]/ [category = "clothes"]/*
The first predicate, [obj_id = "retail"], is evaluated against all objects under the root object, /. Then, the next predicate, [name = "Sales"], is evaluated against the children of the retail database object, which has an object id equal to "retail". The rest of the path expression is evaluated in the same manner as [region = "Asia"] is evaluated against the children of the Sales table object, etc.
Updates: TreeCat can be updated via commit() function, which takes the modifications in the form of a write set. A write set is a collection of (path, value, type) tuples, where path specifies the object to modify, value contains the updated value, and type specifies the operation type, as listed below:
  1. (1)
    Add: Add operation adds a new object of the given path and initializes it with the given value. The operation has a precondition that the parent object exists, but an object with the same path does not.
  2. (2)
    Update: Update operation updates the object of the given path with the given value. If the object does not exist, a new object with the provided value is created.
  3. (3)
    Remove: Remove operation removes the object of the given path and all of its descendants.
  4. (4)
    Merge: Also known as the commit-time update operation (Huang et al., 2020), merge operation turns a read-modify-write operation into a single blind write operation that is performed at commit time. The given object value is a delta, rather than the new object value, and is applied to the existing object value at commit time. The operation has a precondition that the object already exists.
We describe the merge operation in more detail and highlight its importance. Merge operation is useful when objects are frequently updated by read-modify-write operations (e.g., counters), which may result in many read-write conflicts. An important use case in TreeCat is applying changes in statistics that propagate up the hierarchy as a result of any changes in the low level objects, which can result in contention on high level statistics objects. We present an example of an update to table statistics that follows data insert:
Original Value: {...size:1487...min:3...}
Delta: {...size:{op:+,val:124}...min:{op:min,val:0}...}
Final Value: {...size:1611...min:0...}
Currently, the types of delta operations TreeCat supports include addition, subtraction, min, and max for a limited number of types.
Catalog API: We show how catalog API can be built on top of TreeCat with an example from the Spark catalog API we implemented for the experimental evaluation. Each operation takes in an optional TreeTxn, which keeps track of the current transaction session. The following is a simplified snippet of $getTable()$ function, which returns the table information, given a database name and table name.
def getTable(db: String, table: String,
  txn: Option[TreeTxn] = None): CatalogTable = {
  val requestBuilder = queryRequestBuilder(
    f"/[obj_id=\"\$db\" and obj_type=\"database\"]"
    + f"/[obj_id=\"\$table\" and obj_type=\"table\"]")
  setTxnId(requestBuilder, txn)
  val queryResponses = catalogStub
    .executeQuery(requestBuilder.build())
The function constructs a query request with appropriate vid or txn_id if txn is defined. Then, it sends it to the catalog server by a $executeQuery()$ RPC (Remote Procedure Call). queryResponses is a stream of buffers filled with BSON documents, which are deserialized to in-memory CatalogTable object.
5. System Architecture
We first outline the high level architecture of TreeCat, then elaborate on the core system components. We discuss the concurrency control mechanism in Section 6 in greater detail.
5.1. High Level Architecture
TreeCat is designed as a single-server database engine that handles incoming requests from remote client systems by the gRPC interface. It is written in approximately 12000 lines of C++ code, excluding any third-party libraries. Key components include:
  1. (1)
    Backend Server: Provides the API described in the previous section and schedules client tasks to the main thread pool.
  2. (2)
    Executor: Contains utilities for creating the execution plan of a read query and executing it.
  3. (3)
    Transaction Manager: Manages the overall concurrency mechanism. It keeps track of ongoing transaction states and launches a validation and commit process of the transaction upon receiving a commit request.
  4. (4)
    Storage: All catalog data is stored in the storage layer, which provides APIs to get an object by path and traverse all children of an object. The underlying storage engine is RocksDB (Platforms, 2022).
Figure 4. High level architecture of TreeCat
5.2. Backend Server
We use gRPC (gRPC Authors, 2024), a Remote Procedure Call (RPC) framework developed by Google, to provide the APIs listed in the previous section. gRPC is language-agnostic and offers portability across different languages. While REST API is another option with good portability, gRPC provides better performance due to its use of pre-compiled serialization and deserialization routines and support for server-side streaming, especially for large data loads (it is not uncommon for reads in this context to return MBs of data).
5.3. Storage Engine
When selecting the underlying storage engine, we considered several key requirements, including (1) optimizations for efficient traversal of hierarchical data, (2) efficient implementation of explicit versioning semantics, (3) fine-grained concurrency control with strong guarantees that can efficiently handle frequent updates to hierarchical data, and (4) embedded storage for low latency.
Numerous options exist for full-fledged database systems, ranging from RDBMS with XML (Shanmugasundaram et al., 2001; Tatarinov et al., 2002; O’Neil et al., 2004) and graph (ten Wolde et al., 2023) extensions, to graph databases (Neo4j, 2024; NebbulaGraph, 2024; HyperGraphDB, 2024; ArangoDB, 2024; JanusGraph, 2024; Dubey et al., 2016; Zhu et al., [n.d.]; De Leo and Boncz, 2021; Fuchs et al., 2022), and document stores (MongoDB, 2024; CouchDB, 2024; ArangoDB, 2024). However, meeting all the requirements, especially (2) explicit versioning and (3) fine-grained concurrency control, proves to be difficult, as these systems are designed and implemented for more general use cases. The concurrency control mechanism is especially problematic as general schemes adopted by these systems are not specifically designed to handle hierarchical data. The graph or document databases also don’t typically support the fine-grained concurrency control needed to support updates to portions of the hierarchical metadata.
After careful consideration, we built TreeCat on top of RocksDB, a write-optimized key-value store that is widely used as the storage engine of many systems. Inspired by file systems built on top of write-optimized key-value stores (Jannen et al., 2015; Shetty et al., 2013; Esmet et al., 2012), we adopt a storage layout where objects are sorted by the full object path, first by the path depth, then lexicographically. This layout allows fast listing of sibling objects by performing a range scan on the common prefix. It has advantages over storing the parent-child relation as a separate set of edges, which incurs the overhead of joining the objects (vertices) with edges during traversals. Although each range scan incurs a seek cost, the expense is amortized since objects typically have multiple children; as we show later, query cost is bounded by the query’s selectivity because most irrelevant objects are pruned early. This layout is also more suited for versioning and transactional updates compared to the scheme where an object’s children are stored together as a single list. With the latter, installing a new object version forces installation of an updated version of the parent’s children list, potentially cascading changes to the root and causing read–write conflicts that severely limit concurrency under serializable isolation. This also imposes a fundamental limitation of Iceberg for supporting multi-table transactions.
Object values are stored in BSON format (Commons, [n.d.]), a binary JSON format that originated from MongoDB, is queriable, and can easily be converted to JSON format. While other efficient JSON-compatible binary formats exist (e.g., PostgreSQL’s JSONB), we chose BSON for its robust library support and rich feature set, which enables faster system development. We use two separate storage systems, InnerObjectStore and LeafObjectStore, which store non-leaf objects and leaf objects respectively. LeafObjectStore stores one object value for each path and the $vid$ range during which the object is visible. Unlike leaf objects, multiple past versions of non-leaf objects are persisted. For non-leaf objects, we adopt the “time-travel” storage scheme (Wu et al., 2017) where the most recent version is stored in the main SnapshotStore and the chain of past versions are stored in a separate DeltaStore, as illustrated in Figure 5. SnapshotStore has the following mapping:
path $\rightarrow$ delta_vid, cur_vid, object value
where $cur\_vid$ is the $vid$ from which the current snapshot is valid, and $delta\_vid$ is the $start\_vid$ of the previous object value. DeltaStore has the following mapping:
start_vid, end_vid, path $\rightarrow$ object value
where the object value is visible between $start\_vid$ and $end\_vid$. In DeltaStore, the objects are sorted by the object path, then by $start\_vid$ in descending order, so past object values can be traversed through a range scan. Retrieving the object value of a non-leaf object involves first checking the SnapshotStore for the current value and traversing the past values in DeltaStore, if necessary. Retrieving past versions incurs the cost of version traversal, but we expect most operations will access the most recent version. LeafObjectStore has the following mapping:
path $\rightarrow$ create_vid, tombstone_vid, object value/primary path
where the object is visible between $create\_vid$ and $tombstone\_vid$. Because leaf objects can have multiple parents, there could be a layer of indirection where the path could map to the primary path, which, in turn, maps to the object value.
Figure 5. Example storage layout of the Sales table from Figure 2 in InnerObjectStore.
5.4. Query Execution
A read query is executed as a sequence of physical operators analogous to the correlated join operator from RDBMSs (also called CROSS APPLY or LATERAL JOIN), which passes each value from the outer query to a user-defined function or a subquery and evaluates it. Unlike other standard join operators, the correlated join operator allows for early pruning of objects in the inner query by only scanning objects that join with objects that satisfy the outer query. We use this mechanism to avoid scanning large number of leaf objects whose ancestors do not satisfy the preceding predicates in the path expression. For example, while executing /[obj_id = "retail"]/[name = "Sales"]/*/*/*, file objects that belong to the Customer table need not be scanned.
Given a path expression of depth $n$, the execution plan is constructed as a single chain of $n$ execution nodes, each with the corresponding predicate, a fixed-sized output buffer, a pointer to its child node, and handles to the object store. The query plan follows a standard batch iterator model where each node invokes $next()$ on its child node, retrieving a batch of context object paths. For each context object path, the executor scans and evaluates the predicate against every child object visible in the $read\_vid$. If the object satisfies the predicate, its path (or its value if the last execution node) is appended to the output buffer. The execution of a simple example query is illustrated in Figure 6. The planner currently applies one optimization rule where the lower and/or upper bounds of the object store iterator are set if the predicate is only satisfied by a range of object ids (e.g. [1 < obj_id < 100]).
Assuming that the underlying data takes the form of a balanced tree $T$ with a constant fan-out $f$ and height $h$, we can compute the cost of a path query $q$ of depth $d$ where $selectivity(q)=(s_{1},s_{2},..,s_{d})$ specifies the list of predicate selectivities. We also assume independent selectivities for simplicity. The two main cost components are the scan cost and seek cost of every range scan. We first compute the total number of scan operations. At $i^{th}$ level of the query evaluation, the number of scanned objects increases by a factor of $s_{i+1}f$:
$n_{scan}$ $=f+s_{1}f^{2}+...+\prod_{i=1}^{d-1}s_{i}f^{d}$
$\leq\frac{s^{d}f^{d+1}}{sf-1}\text{ where }s=\max_{i=1}^{d-1}s_{i}$
$\leq s^{d}f^{d+1}\text{ when }sf\geq 2$
The number of seek operations is equal to the number of scan operations of the preceding path sub-expression and can be computed to be $s^{d-1}f^{d}$ in a similar manner. Because there is no closed form formula for $f$ in terms of $|T|$, we approximate $f$ to $|T|^{1/h}$. Finally, the total cost can be approximated to $s^{d-1}\cdot|T|^{d/h}(s\cdot|T|^{1/h}\cdot c_{scan}+c_{seek})$ where $c_{scan}$ and $c_{seek}$ represent the cost of a single scan operation and that of a seek operation, respectively. We observe that the multiplier effect of selectivity $s$ effectively bounds the number of scan and seek operations necessary for executing the given query, which aligns with the intuition about “early pruning” of objects. This contrasts with more popular join operators in the relational model, such as sort-merge join or hash join that require at least one scan of the entire table, which can be costly for a large table size (analogously tree size).
5.5. Version Control Operations
The system supports globally consistent chronological versioning of the entire catalog data, using monotone increasing $vid$ associated with every transaction. $snapshot()$ operation is implemented by maintaining an additional mapping between the snapshot name and the $vid$, which can be used to retrieve the correct version. $clone()$ operation copies the source object and all of its descendants (except for the leaf objects) at some $vid$ to the destination path. For cloned leaf objects, the system adds a mapping from the new destination paths to the primary paths, so they are shared between the source and destination parents, saving the storage costs of duplication.
6. Concurrency Control
In this section, we present TreeCat’s concurrency control mechanism, which guarantees serializability. Besides our use case in TreeCat, the scheme is also of independent interest in a more general context of databases with a hierarchical data model.
6.1. Discussion
We first discuss the background and primary motivations, followed by an in-depth description of the implementation details. We adopt the MVOCC scheme where read-only transactions operate on a versioned snapshot of the data, and read-write transactions follow a protocol similar to the standard OCC mechanism. Because we implemented MVOCC on top of the existing version system, $vid$ is synonymous with $timestamp$ as commonly discussed in the concurrency control literature. For read-only transactions, we follow the common practice of assigning a $read\_vid$, the version up to which data are safe for read, at the start of the transaction, and executing queries against the versions of objects visible at the assigned $read\_vid$. Since this aspect of the mechanism aligns with other existing systems, our main focus is on the OCC, which utilizes a predicate-based method, often overlooked by standard schemes.
Figure 6. Query Execution of a Simple Example Query: /[obj_id = retail]/[obj_id = sales]/*. The output buffer in the top most node holds the final result set.
The important feature of TreeCat is that read queries involve predicate read operations. Unlike simple key-value read and write operations, predicate read operations can introduce predicate dependencies (Clark et al., 2024), which are distinguished from the item dependencies (write-read, write-write, and read-write) that are commonly discussed in the concurrency control literature. The infamous phantom read anomaly is one example of a predicate anomaly that can occur if predicate dependencies are not handled correctly by the concurrency control mechanism (Eswaran et al., 1976). Unlike item dependencies—which are based on conflicting operations on the same item—predicate dependencies occur when an update alters whether an object satisfies a given predicate. Specifically, if a write operation’s before-image satisfies the predicate of a predicate read operation but its after-image does not, or vice versa, a predicate dependency exists between the two operations (Adya et al., 2000). Item-based approaches, including record locks in 2PL (2 phase locking), read and write sets in OCC, and time stamps on individual items in TO (timestamp ordering), lack semantics for predicate reads, and cannot handle predicate dependencies without additional mechanisms.
The most common way to augment these schemes to handle predicate dependencies is to lock the entire scan range of the predicate read operation. If an index scan is used, the scan range can be limited to a range of values of the indexed attribute on which the predicate is defined. However, this optimization cannot be applied to arbitrary predicates, which may be a complex expression or involve attributes on which an index does not exist. The scheme is, at best, a coarse-grained approximation that may lock an irrelevant range of values, resulting in unnecessary conflicts.
We illustrate this point with a simple example. Let $R$ be a logical table that is partitioned into multiple files. Transaction $T_{1}$ performs a predicate read operation on $R$’s file metadata with the predicate $p$ on the file statistics. Because there is no index on the file statistics, $T_{1}$ scans the entire file metadata of $R$. Meanwhile, $T_{2}$ inserts a file $f$ with file metadata that does not satisfy $p$ into $R$. There is no reason for the two operations to conflict, because the new file $f$ is never visible to $T_{1}$’s predicate read, regardless of the order of operations. However, with scan range locking, $T_{1}$ would lock the entire range of $R$’s file metadata. In a pessimistic scheme (S2PL), either of the two transactions would block the other. In an optimistic scheme (OCC), if $T_{2}$ is submitted for validation before $T_{1}$, $T_{1}$ would be aborted.
There are predicate-based methods that can precisely detect any predicate dependency. One scheme is to re-execute the scan operations to observe any changes in the visible objects (Larson et al., 2011). While precise, re-executing scans can be costly for large ranges. Precision locking (Jordan et al., 1981) is a form of predicate locking that evaluates the predicate directly against posted write operations. This technique detects predicate dependencies accurately and avoids the satisfiability test problem of conventional predicate locking. But in TreeCat, a read query is correlated whence an object satisfies the entire path expression only if all of its ancestors also satisfy the corresponding path sub-expressions. Hence, a naive implementation that only considers the predicates in a given query can, in fact, result in false conflicts.
In our scheme, we combine scan range locking with precision locking, leveraging the strengths of both methods. We first apply optimistic prefix-based scan range locking to detect candidate conflicts, i.e., write operations that fall within the scan range, but may not necessarily cause conflicts. Then, we use optimistic precision locking to confirm whether the candidate is actually a conflict. This scheme achieves the granularity of precision locking while improving efficiency by pruning unnecessary predicate evaluations with scan range locking. In addition, we make use of an in-memory index structure to quickly locate only the log records that fall in the scan range, rather than applying precision locking on the entire set of log records.
Algorithm 1 Validation Algorithm
1:  for $(path,predicate)\in txn.ScanSet$ do
2:     // scan range locking
3:     if $VersionMap[path]>txn.read\_vid$ then
4:        $iter\leftarrow LogIndex.newIterator()$
5:        $iter.seek(path,txn.read\_vid)$
6:        while $iter.path.parent=path$ do
7:           if $iter.type=merge$ and $iter.image.empty()$ then
8:              $iter.image.construct()$
9:           end if
10:           // precision locking
11:           if $iter.image.before\models predicate$ then
12:              $abort(txn)$
13:           else if $iter.image.after\models predicate$ then
14:              $abort(txn)$
15:           end if
16:        end while
17:     end if
18:  end for
6.2. Implementation
The concurrency control is handled by the TransactionManager, which keeps track of $read\_vid$, the latest version up to which, the catalog can be read safely, and $commit\_vid$, which is the monotone increasing vid assigned to a transaction before final commit. Key in-memory data structures include:
  1. (1)
    $ScanSet$: Maintained by each transaction. Stores pairs of the scan range (in the form of the context object path) and the predicate of the predicate read operations executed as part of the transaction.
  2. (2)
    $LogBuffer$: Maintained by each transaction. Stores before- and after-images of all recent write operations.
  3. (3)
    $VersionMap$: Used by TransactionManager for validation. Maps object path to the most recent $vid$ of its children.
  4. (4)
    $LogIndex$: Used by TransactionManager for validation. Index on the $LogBuffer$, mapping pairs of object path and $vid$ to the log records. The index is implemented using a lock-free skip list and is sorted by the parent path, then by the $vid$.
When $startTransaction()$ is called, a new transaction session is created and $read\_vid$ is loaded as the vid at which all read queries are executed in the transaction session. During query execution, pairs of predicates and parent paths (prefix) of objects scanned for predicate evaluation are added to the $ScanSet$. Also, a private copy of the write set is constructed on the client side. When the transaction is submitted for the commit process, the write set is sent to the server and undergoes preprocessing phase, where it is copied to the $LogBuffer$ and preconditions are checked. If the transaction passes these checks, the transaction is scheduled for the main commit process.
The main commit process consists of two phases: validation and batch write. For each phase, there is a dedicated thread pool of a fixed size, with each thread handling a single hash partition of the $ScanSet$ and the write set. At the start of the validation, TransactionManager executes a $fetch\_and\_add()$ operation on the $commit\_vid$ and assigns it to the transaction. Then, each worker thread validates the scan ranges in the $ScanSet$ against the $VersionMap$ to determine whether any objects scanned by predicate read operations have changed. If potential conflicts are detected, an iterator on $LogIndex$ is set to scan and apply precision locks on corresponding recent writes with $vid$ greater than the transaction’s $read\_vid$ and fall in the scan range. If any conflict is detected through precision locking, the transaction is aborted. After all threads pass validation, they update the $VersionMap$ and $LogIndex$ based on their respective write set partitions for the transaction. The validation phase (omitting the initial $vid$ assignment and latter updates to in-memory data structures) proceeds as per Algorithm  1. Note that for merge operation, before- and after-image may have to be constructed as they are not available until the final commit. However, once constructed, the images are cached so they are accessible during the validation of subsequent transactions.
Once validation is complete, the transaction proceeds to the write phase. During this phase, each thread processes its assigned write partition, evaluating the final values of merge operations (commit-time updates) and applying the writes as an atomic batch write operation to the underlying RocksDB instance. After every thread finishes its batch write, WAL records of the transaction are flushed to disk along with those of other transactions via $SyncWAL()$ operation, which calls $fsync()$ on the WAL files. The $read\_vid$ is also updated, making the committed changes visible for subsequent reads. Transactions are processed one at a time in each phase, but parallelism is achieved via hash-partitioning the workload and pipelining different phases of the commit process.
Old $LogBuffer$s and entries in $VersionMap$ and $LogIndex$ that are no longer needed are asynchronously garbage collected by a background thread. For this purpose, the TransactionManager keeps track of $watermark$, the oldest $read\_vid$ among the current transactions. After each transaction commit, the $watermark$ is updated and all entries in the aforementioned data structures with $vid$ below the $watermark$ are garbage collected.
6.3. Proof of Correctness
In this section, we prove that TreeCat’s concurrency control mechanism guarantees serializable isolation.
Concurrency Control Model: We first formalize the concurrency control model. We follow Adya’s model (Adya et al., 2000), where every object in the database has one or more versions. Every operation on an object is mapped to exactly one version of that object. Given an object $x$, $x_{i}$ is a version with $vid$ $i$, installed by transaction $T_{i}$. Every object is assumed to exist forever regardless of when they are inserted or deleted. The unborn state is represented as version $x_{init}$, while the dead state is represented as $x_{i}$, installed by the corresponding delete operation of $T_{i}$. Every object is uniquely identified by its full path.
It is important to discuss how TreeCat’s query evaluation is represented in this model. Evaluation of a path expression involves performing a series of predicate range scans, each with respect to a context parent object. For example, given the path expression, /[obj_id = retail]/[name = "Sales"], the evaluation of the second predicate, $name=``Sales"$, is executed as a range scan on the children of /retail. Each range scan can be modeled as a predicate read operation, where the predicate is a conjunction of the original predicate and the equivalence between the object parent path and the context parent path. Following the above example, evaluating $name=``Sales"$ against /retail’s children objects can be modeled as a predicate read operation with the predicate, $name=``Sales"\wedge parent=/retail$. It should be noted that the evaluation of a single predicate in a path query can correspond to multiple such predicate read operations, as there can be multiple context parent objects with which the predicate is evaluated.
Main Proof: We now proceed with the main proof.
Lemma 1. For any two committed transactions $T_{i}$ and $T_{j}$ where $commit\_vid_{i}<commit\_vid_{j}$, there is no dependency from $T_{j}$ to $T_{i}$.
Proof.
We prove the lemma by enumeration. First, there cannot be any ww-dependency from $T_{j}$ to $T_{i}$ because write operations are performed in the order of the $commit\_vid$ in the write phase. Second, wr-dependency from $T_{j}$ to $T_{i}$ is also not possible. Given that $read\_vid_{i}<commit\_vid_{i}$ and $commit\_vid_{i}<commit\_vid_{j}$, $read\_vid_{i}<commit\_vid_{j}$ by transitivity. Because changes committed by $T_{j}$ become visible only after the $read\_vid$ has been incremented to $commit\_vid_{j}$, $T_{i}$, which has a lower $read\_vid$, cannot observe any changes installed by $T_{j}$. Lastly, we show by contradiction that there cannot be any rw-antidependency from $T_{j}$ to $T_{i}$. Assume that there is a rw-antidependency from $T_{j}$ to $T_{i}$. Consider the state when $T_{i}$ passed validation, but $T_{j}$ is about to enter validation. There must be a write operation $w$ executed by $T_{i}$ on an object $o$, whose before- or after-image satisfies the predicate of a predicate read operation $P$ executed by $T_{j}$, which scans $w$’s before-image. Per our earlier discussion, the predicate of $P$ is $p\wedge parent=c$, where $p$ is the predicate in the path query and $c$ is the context parent object. During the query execution of $T_{j}$, $(c,p)$ was added to $T_{j}$’s $ScanSet$. Also, $(o.parent,i)$ was added to $VersionMap$ and a mapping from $o$ to $w$’s before- and after-images was added to $LogIndex$ during the validation of $T_{i}$. During the validation of $T_{j}$, $c$ is found in $VersionMap$ because $c=o.parent$ holds true. Since $read\_vid_{j}<commit\_vid_{i}$, $p$ is evaluated against $o$’s entries in $LogIndex$. Because the before- or after-image of $w$ satisfies $p$, the rw-conflict is detected and $T_{j}$ is aborted, leading to a contradiction. We have now considered all types of dependencies, so the lemma holds true. ∎
Theorem 1. Any schedule of committed transactions output by TreeCat is conflict serializable.
Proof.
We prove the theorem by contradiction. Assume that there is a schedule $s$, output by TreeCat, that is not conflict serializable. Then, the corresponding precedence graph, $G(s)=(V,E)$ must contain at least one cycle. Consider a directed edge $(T_{i},T_{j})$ in the cycle. Because there is a dependency from $T_{i}$ to $T_{j}$, $commit\_vid_{i}<commit\_vid_{j}$ by Lemma 1. Because $(T_{i},T_{j})$ is in a cycle there is a path from $T_{j}$ to $T_{i}$. Using Lemma 1 and transitivity, we can prove by induction (which we omit for brevity) that $commit\_vid_{i}>commit\_vid_{j}$. We end up with a contradiction, so the theorem holds true. ∎
System MVCC Optimistic Pessimistic Multi-table Transaction
AWS Redshift
Google BigQuery
Snowflake
Vertica
HMS
Delta Lake
Iceberg Nessie Catalog
Hudi
TreeCat
Table 1. Concurrency control mechanisms of OLAP systems.
Figure 7. Throughput and abort rate across different number of client threads and different types of workloads.
6.4. Related Work
We first discuss the concurrency control mechanisms of other catalog systems and then examine how predicate reads are handled by concurrency control mechanisms across more general data systems.
Catalog Concurrency Control: Traditional RDBMSs store catalog as a set of relations, relying on the same concurrency mechanism that is used for data. While data warehouses use proprietary mechanisms, Snowflake is known to use FoundationDB to persist transaction states and lock queues (Motivala, 2018). Even though resource locks at the level of micro-partitions are available, most single table DML queries require locking all micro-partitions (Snowflake, 2023). Delta Lake uses snapshot-based OCC where, during the final commit, it tries to atomically write the next delta log record. If the underlying storage system does not support atomic file rename, it has to resort to another transactional system, such as DynamoDB (Scott Sandre and Kryński, 2022). Delta Lake currently only supports single-table transactions. However, there is early work on introducing a new commit protocol called commit coordinator, an external system that is dedicated to handling concurrency control, which may support multi-table transactions in the future (Tathagata Das et al., 2024). Iceberg also uses snapshot-based OCC where, after writing all the metadata files (manifest files, manifest list files, and metadata file), it does a final commit by switching the current pointer in the catalog to the new metadata file. If there is a conflict, the transaction undergoes a retry validation, so the entire operation need not be repeated from scratch (Okolnychyi et al., 2024). Like Delta Lake, Iceberg’s commit protocol only supports single-table transactions. Recently, multi-table transactions have been added to the Iceberg REST API. But as far as we know, Nessie (Nessie, [n.d.]), which supports git-like versioning, is the only compatible catalog that supports it. Even so, the transactions operate at the level of a single branch of the whole database. Lastly, Hudi also uses snapshot-based OCC and does conflict resolution based on the files modified by multiple writers (Foundation, 2024b). Hudi requires an external lock provider for locking tables and can only support single-table transactions.
Predicate Read: In order for a data system that supports predicate read operations to guarantee serializable isolation, it has to account for any predicate dependencies in addition to item dependencies (Adya et al., 2000). The standard solution is to lock the entire scan range of the predicate read operation, an example of which is next key locking (Mohan et al., 1989), which locks the current record and the gap until the next record. Since taking many record locks can cause excessive overheads, multiple granularity locking (Gray et al., 1976) is commonly used to escalate locks to more coarse-grained ones, such as table locks. Large scan range may cause unnecessary conflicts, so it is crucial to minimize the scan range. For simple predicates, performing an index scan can limit the scan range. A more advanced technique may be used to better approximate the minimum scan range for complex predicates. For example, Spanner uses a runtime data structure called a filter tree to compute the approximate scan range (Bacon et al., 2017). One way to precisely handle predicate dependencies is to validate the scan operation by re-executing it, which is adopted by Hekaton (Diaconu et al., 2013). Another approach is to use predicate-based locking. Conventional predicate locking involves satisfiability test problem, which is known to be NP-complete (Hunt and Rosenkrantz, 1979). However, precision locking avoids this problem by directly evaluating predicate against before- and after-images of posted writes (Jordan et al., 1981). The technique was not used in any real system for a long time, but was later adopted by Neumann et al. for implementing the MVOCC mechanism of Hyper (Neumann et al., 2015).
7. Experimental Evaluation
In this section, we present a comprehensive experimental evaluation of TreeCat. We set up a cluster of 4 physical machines on CloudLab (CloudLab, 2025), each equipped with a 10 core Intel Xeon Silver 4114 processor, 192GiB DDR4 memory, and a 64GB SATA SSD, interconnected by 10Gbps ethernet. We implemented a catalog API, including $getTable()$ and $listFiles()$ operations, in Apache Spark 3.4, allowing communication with a standalone instance of TreeCat.
7.1. Concurrency Control Mechanism
To evaluate the concurrency control mechanism, we implemented two standard concurrency control protocols in TreeCat. The first is OCC with optimistic scan range locking (OSL), which also primarily utilizes prefix-based scan range locks. However, instead of precision locking, it uses more fine-grained range locks for bounded range scans. For example, if only table partitions 1 to 100 are scanned, only the range of object id, $[1,100]$, rather than the entire range of table partitions, is optimistically locked. The second scheme implements S2PL with multiple granularity locking (MGL). All three schemes efficiently handle read-only transactions with an identical MVCC protocol. Our new proposed concurrency control mechanism, i.e., TreeCat’s default protocol, is referred to as OSPL (optimistic scan range and precision locking).
We used a microbenchmark largely derived from the TPC-DS benchmark. While TPC-DS proceeds as a single fixed sequence of phases, we design our benchmark so that multiple client threads randomly select queries and perform the corresponding metadata operations (against TreeCat) concurrently. Besides the original read queries and data maintenance operations, which involve data insertion and deletion on the fact tables, we introduce two new operations: an optimize operation, which selects and merges small files, and dimension table insertion. These new operations allow the simulation of real-world scenarios that require support for multi-table transactions. The relative distribution of write operations is adjusted with realistic assumptions about standard data warehouse workloads. For example, we assume that fact table insertion occur 12 times more frequently than dimension table insertion (every 5 minutes vs. every hour) and 144 times more frequently than optimize operations (every 5 minutes vs. every 12 hours). We adopted the schema of the TPC-DS benchmark, but used a custom data generator to generate relatively large amount of metadata. The amount of metadata per table, the cardinality of surrogate keys and business ids are in line with the TPC-DS specification for a scale factor of 100 terabytes. In our experiment, we deployed TreeCat on a single machine and client threads on the other three machines.
Figure 8. 99th percentile operation latency of read-write transactions across different number of client threads and operation types for 30 threads, respectively. In the latter figure, we use abbreviations: Opt (optimize), Fact (fact table insertion), Dim (dimension table insertion), and Del (delete).
We first conducted a scalability experiment to measure overall throughput, abort rate, and latency over a set time duration across different numbers of client threads and percentages of read-write transactions. For all three schemes, read-only transactions do not cause any conflict as they follow the MVCC protocol. Consequently, performance differences between the schemes are minimal for more read-intensive workloads, as shown in Figure 7. However, for write-intensive workloads, throughput diverges, with OSPL significantly outperforming the other two as the number of client threads increases. The main source of contention is read-write conflicts on the customers table, which is one of the dimension tables to which data is inserted periodically. Inserting into a fact table requires joining the input data with dimension tables on the business ids (primary key in the source OLTP database) and should only conflict with dimension table insertion if the ranges of dimension table business IDs overlap. However, both OSL and MGL apply coarse-grained scan range lock on the dimension tables, leading to unnecessary conflicts. As a result, the abort rate of OSL increases almost linearly with the number of threads, reaching 31.3% for the balanced workload and 63.7% for the write-intensive workload at 30 client threads. Similarly, MGL maintains low throughput since transactions are blocked while waiting for lock acquisition. By contrast, OSPL avoids false conflicts through precision locking on the file statistics, achieving significantly higher throughput and a low abort rate. The latter two plots of Figure 7 show this trend more clearly as the percentage of read-write transactions increases for 30 client threads. A considerable proportion of read-only transactions scan entire fact tables and have longer latency than read-write transactions, which is why the throughput of OSPL increases with more write-intensive workloads.
Next, we compare latency across the different schemes for the balanced workload. Because read-only operations have longer latencies than read-write transactions and follow the same MVCC protocol in all three schemes, we isolate and analyze the latencies of read-write transactions. Figure 8 shows the 99th percentile latency of read-write transactions across different number of threads and for each operation type. While the latency remains consistent for the two OCC schemes, it increases almost linearly with the number of client threads for MGL as transaction lock queues build up. Categorizing transactions by operation type reveals that the latency of dimension table insertion is almost as twice as long as that of fact table insertion, indicating that the conflicts on the updated dimension table is the main source of contention for MGL. The latency of optimize operation is significantly longer for OSPL compared to OSL because higher number of small files accumulate with less aborts on insertion, resulting in larger optimize operation.
Figure 9. Throughput for read-intensive (80-20) and write-intensive (20-80) workloads with multiple databases (10 client threads per database), respectively.
Lastly, we evaluate how throughput scales across different numbers of databases for two types of workload with low contention (10 threads per database). As expected, all three schemes have very similar throughput for the read-intensive workload, saturating around 3 databases. For the write-intensive workload, OSPL generally outperforms the others. However, the throughput converges with that of MGL as it is increasingly constrained by contention on physical resources. OSPL still outperforms MGL at 6 databases by a small margin. OSL exhibits significantly lower throughput, compared to the other schemes, and also reaches the saturation point earlier at 3 databases due to high abort rate.
7.2. Comparative Evaluation
Next, we compare TreeCat with HMS, Delta Lake, and Iceberg. An apples-to-apples comparison is difficult because TreeCat does not support a catalog API fully compatible with Spark and all the systems have different architectures and transactional guarantees. We nonetheless try our best to compare the relative costs of metadata operations. We set up HDFS (Hadoop Distributed File System) and Spark on the 4 node cluster. We used Hive 2.3, Delta Lake 2.4, and Iceberg 1.5, all of which are compatible with Spark 3.4. We used a single table with a schema identical to that of the Store Sales table in the TPC-DS, partitioned by the sales date, from 1998-01-01 to 2003-12-31. To test scalability of read performance, we populated the table with varying numbers of files and measured the latency of retrieving files with a range filter on the date. While our primary implementation of the catalog API uses a single thread to retrieve the query result via a single stream, both HMS and Iceberg leverage multiple cores on the client machine to retrieve the result over multiple streams where HMS calls $listFile()$ and Iceberg retrieves manifest files in parallel. Client-side parallelism can be achieved with TreeCat quite easily by adding a partitioning predicate to the path query, so that multiple threads can concurrently submit queries and retrieve distinct result sets. Although a hash function combined with the $mod$ operation would typically be the preferred approach, our implementation applies the $endwith()$ operation to the partition value due to the limitations of the current query language. We used this parallel implementation for the following experiments.
Figure 10. Median latency of file retrieval across different number of files for 1 day and 365 days predicate ranges, respectively.
We first analyze the scalability of the file retrieval operation with high selectivity (date range of 1 day) and with low selectivity (date range of 365) across different numbers of files. For the 1-day predicate range, TreeCat consistently shows low latency as it efficiently traverses the hierarchical data on the server and only returns the file metadata in the given data partition. HMS also scales quite well as it calls $listPartitionsbyFilter()$ to only retrieve the partitions in the given range from the thrift server. Although Delta Lake benefits from Spark’s distributed processing engine, the overhead of setting up and executing the job adds significant latency, resulting in a relatively high minimum latency of around 600 ms. This contrasts with previous experimental results (Armbrust et al., 2020; Jain et al., 2023) and we suspect that performance could be significantly improved with a better-configured, larger Spark cluster. But the client-side implementation of getting the end result appears to be a major bottleneck. Although Iceberg utilizes multiple cores for processing manifest files, the size of an individual manifest file can get large, which contributes to the longer latency for the higher file counts. TreeCat outperforms the other systems and scales well also for the 365-days predicate range. HMS beats both Delta Lake and Iceberg when the file count reaches 50000, which we mainly attribute to the relatively low latency of HDFS, compared to object stores.
Figure 11. Median latency of file retrieval across different predicate ranges, and numbers of physical cores on the client machine, respectively. File count is 50000 for both while predicate range is 365 for the latter.
Next, we analyze how the selectivity of the date range filter affects the latency for the file count of 500000. Both TreeCat and HMS effectively prune irrelevant date partitions, resulting in better performance for narrower predicate ranges. Iceberg, on the other hand, operates at the granularity of manifest files, so the latency remains relatively high. Delta Lake also does not scale too well as the number of retrieved files increases. We also evaluate how the number of physical cores on the client machine affects the overall performance, using 365-day date range and a file count of 500000. As expected, the performance of HMS, Iceberg, and TreeCat, which take advantage of client-side parallelism, degrades with limited number of cores while that of Delta Lake remains consistent.
Figure 12. Throughput and latencies of $alterTable()$.
Lastly, we measure the relative cost of the commit protocol for each system by invoking a light-weight metadata operation, $alterTable()$, on an empty Store Sales table. The experiment is only intended to provide a rough estimate as the systems have different client-side implementations and provide different transactional guarantees (which HMS does not even provide). We measure both the total throughput over a set duration and end-to-end latency of each invocation. We observe that lakehouse storage formats, especially Delta Lake, have significantly longer latencies with median latency of 1434.1 ms for Delta Lake and 180.4 ms for Iceberg (throughput of 0.6 ops/s and 5.6 ops/s, respectively). Delta Lake and Iceberg first have to reconstruct the most current snapshot of the table, which tends to be costly. In addition, the final commit protocol involves creating metadata files and calling an atomic operation on the underlying HDFS, which contributes to the extra latency. As a transactional system, TreeCat exhibits significantly lower median latency of 8.4 ms and higher throughput of 117 ops/s.
8. Conclusion
Motivated by the increasing use of a diverse set of data management systems within a single enterprise and the rise of disaggregated architectures, in this paper, we identified the critical systemic and functional requirements for a standalone catalog and justified the key design decisions for our catalog engine, TreeCat. By adopting a hierarchical data model, introducing a specialized data format and layout, and developing a novel concurrency control mechanism, we directly address emerging use cases that challenge state-of-the-art operational metadata systems. Our experimental evaluation not only exposes the limitations of existing solutions but also validates the advantages of our design choices. Looking ahead, we are actively pursuing further enhancements, including (1) addressing additional requirements outlined in Section 3, (2) implementing and testing more complex workloads—such as processing non-tabular data, provenance, and versioning, and (3) developing a comprehensive catalog API to enable complete end-to-end evaluation.